OpenStudy (anonymous):

Help Please! I seriously don't understand this. Thanks. Verify each identity. 2. sin(x+π)/cos(x+3π/2) = tan^2 x - sec^2 x

OpenStudy (nincompoop):



OpenStudy (nincompoop):



OpenStudy (anonymous):

can you keep on explaining it, like I have the sheet with all the identities and pellet, I just seriously don't understand how to do them. Like I have 24 questions, and I've understood legit 5 @nincompoop

OpenStudy (nincompoop):

did you look at the cheat sheet link I provided?

OpenStudy (anonymous):

yup, and I feels like I'm looking at Chinese. like honestly i just completely suck at trig, i had an a in pre-cal until this, and like my teacher sucks, and the book does not explain at all

OpenStudy (nincompoop):

books do explain, it is a matter of comprehension

OpenStudy (nincompoop):

look at the second page of the link

OpenStudy (anonymous):

like no i teach myself everything from the book before a test and i ace it, cause im a self-learner and i seriously cannot focus in a class but this particular subject i just cant seem to teach myself

OpenStudy (anonymous):

yes, i did and like i said. chinese. like i think you expand the numerator and the denominator with the Sum and Difference Formulas, but then im just lost

OpenStudy (nincompoop):

eliminate the cosine on the left hand side by multiplying both sides by it

OpenStudy (nincompoop):

express the right-hand side by sin and cos

OpenStudy (anonymous):

i dont understand that at all...

OpenStudy (nincompoop):

what will the expansion of numerator and denominator achieve? your steps must have some purpose to what you're trying to achieve.

OpenStudy (nincompoop):

eh did you learn trigonometry in algebra lessons?

OpenStudy (anonymous):

im trying to prove the identity

OpenStudy (nincompoop):

yeah we can prove the equation to be true

OpenStudy (nincompoop):

if you followed the two steps I provided then write here what you have. we will proceed with the rest of the steps

OpenStudy (anonymous):

Left side of the equation: sin (x + Pi) = -sin x ; and cos (x + 3Pi/2) = sin x sin ( x + Pi)/cos (x + 3Pi/2) = -sin x/sin x = -1(1) Right side: tan^2x - 1/sin^2x = sin^2x/cos^2 x - 1/cos^2x = (sin^2 x - 1)/cos^2x = = -(1 - sin^2x)/cos^2 x= -cos^2 x/cos^2x = -1 (2) Finally, compare left side (1) to right side (2).

OpenStudy (anonymous):

not really, i mean i took algebra in 7th grade so i dont even remember anything. i just dont understand how to apply what you're saying. can you like draw it so i can visually see it

OpenStudy (anonymous):

Correction: line 5: Right side: tan^2 x - 1/cos^2 x = sin^2 x/cos^2 x - 1/cos^2 x =...

OpenStudy (anonymous):

wait, so for the left side shouldn't it like equal the right side and not -1 cause your like proving it to be true?

OpenStudy (anonymous):

The left side equals the right side. This proves the identity.

OpenStudy (anonymous):

oh, so i actually understood what you did @thu1935, but like im supposed to do it differently you see i actually attached the correct answer, cause my teacher told us the answers. but, thank you so much for your help anyway, you were a lot more helpful than the other person who answered.

OpenStudy (nincompoop):

I doubt that you learned anything, but good luck anyway